# Knowledge-guided machine learning for county-level corn yield prediction under drought

### v2

@@ 5.1 Spatial influence of soil moisture on model prediction
To answer the first question, we visualized the attention scores of soil moisture across twelve states in the U.S. Corn Belt from June to August over the years 2019 to 2023 (Fig. 8). The attention scores indicated the relative importance assigned to soil moisture by the model in different regions, with higher scores suggesting a stronger influence on yield prediction. The attention map highlighted how the model’s reliance on soil moisture varied across different growth stages and drought conditions, allowing us to assess whether soil moisture has a greater impact on the model in drought-affected areas. To enhance visualization, we normalized attention values within each year. Consequently, analysis should focus on attention trends across regions within the same year rather than comparing values across different years.
@@
In June, during the early growth period, the attention scores for soil moisture were generally lower across all five years, This might be because the temperature was not too high and drought conditions were not severe at this time, leading to a less pronounced correlation between soil moisture and corn yield. In July, attention scores increased, particularly in drought-affected areas, as soil moisture became more influential during active growth and vegetative stages. By August, attention scores peaked in drought-affected states, aligning with the critical reproductive phase of corn when adequate soil moisture was essential for kernel development.
From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations, aligning with the drought severity in different years. In early 2019 (Fig. 8(a1)), attention concentrated in Iowa. Subsequently (Fig. 8(a3)), regions with higher attention values expanded gradually, though Iowa retained the highest attention, aligning with observed drought conditions that year. During 2020 (Fig. 8(b)), high-attention regions were initially concentrated in Iowa and Nebraska, later expanding to adjacent areas. By 2021 (Fig. 8(c)), attention intensified over North Dakota, South Dakota, Minnesota, and Nebraska. In 2022 (Fig. 8(d)), high-attention areas were primarily located in Nebraska, Kansas, South Dakota, and Iowa. In 2023 (Fig. 8(e1)), attention initially focused on Kansas, expanding thereafter across a majority of states. Ultimately, significant attention emerged in Nebraska, Kansas, Minnesota, Iowa, Wisconsin, and Missouri, indicative of substantial drought conditions (Fig. 8(e3)). Overall, the period from 2019 to 2023 exhibited a trend of drought conditions and corresponding attention expanding towards the central United States, consistent with actual observations. The reason for the incomplete consistency between the two could be that although drought was present in the region, it did not severely impact corn yield. Alternatively, even without drought in a particular area, other factors, such as the derecho storm, might have affected corn yield. This analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions.
@@ 5.2 Temporal role of soil moisture during the corn growth season
To answer the second question and illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—every 16 days from June to September in drought-affected states (Fig. 9). We found that the VIs had the highest influence on the model around August, which aligned with findings from previous study (Johnson, 2014). In this period, VIs showed the strongest correlation with corn yield because this period corresponded to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield (Huang et al., 2014). High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation (Wang et al., 2005). Additionally, we found that weather data showed a significant increase in attention around July in 2020 (Fig. 9(b)) and 2022 (Fig. 9(d)), and the periods with noticeable attention spikes aligned with the trends of VIs. This phenomenon was more pronounced in soil moisture, as soil moisture attention showed a strong correlation with VIs in all years except 2022. In 2022, a slight increase in soil moisture attention could still be observed near the VIs peak. This might be because VIs during these periods are closely related to certain weather data and soil moisture data. For example, NDWI increases with higher PPT because more rainfall enhances soil moisture and plant water content (Ashok et al., 2021). This indicated that our attention mechanism effectively captured feature importance over time dimension.
@@ 5.3 Statistical impact of soil moisture in drought and non-drought regions
To answer the third question, The box plot was used to illustrate the distribution of soil moisture attention across all counties from 2019 to 2023 (Fig. 10). Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought-affected areas and blue-marked non-drought areas from 2019 to 2023 based on Table 7. Across all five years of attention values, drought-affected areas exhibited fewer outliers compared to non-drought areas, suggesting that the model’s attention to soil moisture is more stable in drought-affected regions. Additionally, in all years, the median of soil moisture attention in drought-affected areas was consistently higher than in non-drought areas, indicating that soil moisture had a greater impact on the model in these areas.
Figure 10: The boxplot of soil moisture attention in drought-affected areas (D--A) and non-drought areas (N--D) across all counties.
@@ 5.4 Interpreting corn yield prediction errors via soil moisture prediction anomalies
To answer the fourth question, we plotted the absolute error map of both soil moisture and corn yield predictions (Fig. 11). In our modeling process, we treated soil moisture as an intermediate variable to guide the corn yield output, so when corn yield prediction was inaccurate, the corresponding soil moisture might also exhibit anomalies. This analysis aims to conduct interpretability analysis of inaccurate corn yield predictions. In 2019 (Fig. 11(a)), the drought-affected state Michigan exhibited relatively accurate corn yield predictions from the KGML-SM model; however, the soil moisture output was less accurate. This suggested that soil moisture was challenging to predict in drought-affected areas. In 2020 (Fig. 11(b)), inaccurate corn yield prediction in Iowa was primarily attributable to the derecho storm’s impact (Hosseini et al., 2020), which soil moisture could not reflect. In 2021 (Fig. 11(c)), drought in North and South Dakota resulted in inaccurate corn yield predictions, and corresponding soil moisture predictions were also found to be inaccurate. In 2022 (Fig. 11(d)), Nebraska, Kansas, and South Dakota experienced drought, and their soil moisture predictions were also inaccurate. In 2023 (Fig. 11(e)), Iowa and Missouri were affected by drought. Iowa’s corn yield and soil moisture prediction were accurate, which indicated that our model performed well. But the prediction for Missouri exhibited errors, despite relatively accurate corresponding soil moisture predictions for the state. According to report (Ward, 2023), some damaging winds and hail significantly impacted corn field, causing root lodging in late June and early July, particularly around Mooresville and Rockport, Missouri. Hail damage was also reported at Garden City, Missouri. These types of mechanical damage were not reflected in standard soil moisture data and constituted yield-limiting factors likely contributing to the corn yield prediction inaccuracies in Missouri. The analysis indicates that a significant portion of the error in corn yield prediction is related to soil moisture prediction error, given the key role soil moisture plays in corn growth; thus, if soil moisture is not accurately predicted, the accumulated error in modeling will lead to inaccurate corn yield prediction.
@@ 6 Conclusion
In this study, we proposed the KGML-SM model, where the W2S encoder is designed to capture the influence of weather on soil moisture, and the attention module is employed to weight different input features for final corn yield prediction. To address the known issue that drought conditions often lead to yield overestimation, we introduced a drought-aware loss function to mitigate this problem in drought-affected regions. We constructed both an APSIM field-level dataset and a GEE county-level dataset, leveraging a process-based model to learn the corn growth process by pretraining the KGML-SM model on the APSIM field-level dataset and finetuning it on the GEE county-level dataset. Our analysis covered 12 states in the U.S. Corn Belt to investigate the impact of soil moisture on corn yield prediction. The proposed method consistently outperformed baseline models across multiple test years. Furthermore, we studied the spatial and temporal influence of soil moisture through attention visualization, revealing when and where the model places greater focus on soil moisture. Finally, based on the relationship between soil moisture and corn yield prediction, we investigated the causes of inaccuracies in corn yield prediction and provided an explanation. In future work, we aim to apply transfer learning techniques to adapt models trained on well-studied regions with abundant simulated data to regions with limited data availability.

### v3

@@ 5.1 Spatial influence of soil moisture on model prediction
To answer the first question, we visualized the attention scores of soil moisture across twelve states in the U.S. Corn Belt from June to August over the years 2019 to 2023 (Fig. 8). The attention scores indicated the relative importance assigned to soil moisture by the model in different regions, with higher scores suggesting a stronger influence on yield prediction. The attention map highlighted how the model’s reliance on soil moisture varied across different growth stages and drought conditions, allowing us to assess whether soil moisture has a greater impact on the model in drought-affected areas. To enhance visualization, we normalized attention values within each year. Consequently, the analysis focused on attention trends across regions within the same year, while cross-year comparisons were not meaningful due to the normalization.
@@
In June, during the early growth period, the attention scores for soil moisture were generally lower across all five years, This might have been because the temperature was not too high and drought conditions were not severe at this time, leading to a less pronounced correlation between soil moisture and corn yield. In July, attention scores increased, particularly in drought-affected areas, as soil moisture became more influential during active growth and vegetative stages. By August, attention scores peaked in drought-affected states, aligning with the critical reproductive phase of corn when adequate soil moisture was essential for kernel development.
From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations. In early 2019 (Fig. 8(a1)), attention concentrated in Iowa, which does not align with the drought and yield reduction observed in North Dakota. An explanation is the widespread prevented planting reported that year, with the highest levels occurring in parts of eastern South Dakota, northwest Ohio, northeast Illinois, southwest Minnesota, and along the Mississippi (American Farm Bureau Federation, 2019; Farm Service Agency, U.S. Department of Agriculture, 2019), which disrupted normal planting schedules and shifted yield risks beyond drought-affected areas. During 2020 (Fig. 8(b)), high-attention regions were initially concentrated in Iowa and Nebraska, later expanding to adjacent areas. By 2021 (Fig. 8(c)), attention intensified over North Dakota, South Dakota, Minnesota, and Nebraska. In 2022 (Fig. 8(d)), high-attention areas were primarily located in Nebraska, Kansas, South Dakota, and Iowa. In 2023 (Fig. 8(e1)), attention initially focused on Kansas, expanding thereafter across a majority of states. Ultimately, significant attention emerged in Nebraska, Kansas, Minnesota, Iowa, Wisconsin, and Missouri, indicative of substantial drought conditions (Fig. 8(e3)). Overall, the period from 2019 to 2023 exhibited a trend of drought conditions and corresponding attention expanding towards the central United States, consistent with actual observations. The reason for the incomplete consistency between the two could be that although drought was present in the region, it did not severely impact corn yield. Alternatively, even without drought in a particular area, other factors, such as the storm or flooding, might have affected corn yield. This analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions.
@@ 5.2 Temporal role of soil moisture during the corn growth season
To answer the second question and illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—at 16-day intervals from June to September in drought-affected states (Fig. 9). We found that the VIs had the highest influence on the model around August, which aligned with findings from previous study (Johnson, 2014). In this period, VIs showed the strongest correlation with corn yield as this period corresponded to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield (Huang et al., 2014). High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation (Wang et al., 2005). Additionally, we found that weather data showed a significant increase in attention around July in 2020 (Fig. 9(b)) and 2022 (Fig. 9(d)), and the periods with noticeable attention spikes aligned with the trends of VIs. This phenomenon was more pronounced in soil moisture, as soil moisture attention showed a strong correlation with VIs in all years except 2022. In 2022, a slight increase in soil moisture attention could still be observed near the VIs peak. This might be because VIs during these periods are closely related to certain weather data and soil moisture data. For example, NDWI increases with higher PPT because more rainfall enhances soil moisture and plant water content (Ashok et al., 2021). This indicated that our attention mechanism effectively captured feature importance over time dimension.
@@ 5.3 Statistical impact of soil moisture in drought and non-drought regions
To answer the third question, the box plot was used to illustrate the distribution of soil moisture attention across all counties from 2019 to 2023 (Fig. 10). Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought-affected areas and blue-marked non-drought areas from 2019 to 2023 based on Table 9. Across all five years, drought-affected areas exhibited fewer outliers compared to non-drought areas, suggesting that the model’s attention to soil moisture is more stable in drought-affected regions. Additionally, in all years, the median of soil moisture attention in drought-affected areas was consistently higher than in non-drought areas, indicating that soil moisture had a greater impact on the model in these areas.
Figure 10: The boxplot of soil moisture attention in drought-affected areas (D-A) and non-drought areas (N-D) across all counties.
@@ 5.4 Interpreting corn yield prediction errors via soil moisture prediction anomalies
To answer the fourth question, we plotted the absolute error map of both soil moisture and corn yield predictions (Fig. 11). In our modeling process, we treated soil moisture as an intermediate variable to guide the corn yield output, so when corn yield prediction was inaccurate, the corresponding soil moisture might also have exhibited anomalies. This analysis aims to interpret inaccurate corn yield predictions.
In 2019 (Fig. 11(a)), the absolute error map shows that corn yield predictions in North Dakota exhibited relatively large errors, while the corresponding soil moisture predictions remained accurate. This inconsistency suggests that yield errors were not primarily driven by soil moisture misrepresentation but were instead likely associated with large-scale prevented planting events in 2019 (American Farm Bureau Federation, 2019; Farm Service Agency, U.S. Department of Agriculture, 2019), which cannot be fully explained by soil moisture dynamics. In 2020 (Fig. 11(b)), inaccurate corn yield prediction in Iowa was primarily attributable to the derecho storm (Hosseini et al., 2020), which soil moisture could not reflect. In 2021 and 2022 (Fig. 11(c–d)), the model exhibited large corn yield prediction errors in drought-affected regions such as North and South Dakota, Nebraska, and Kansas. These yield errors coincided with inaccurate soil moisture predictions in the same areas, suggesting that misrepresentation of soil moisture dynamics under drought conditions was a key contributor to yield prediction inaccuracies. In 2023 (Fig. 11(e)), prediction errors were observed in Missouri and Minnesota. In Missouri, however, the corresponding soil moisture predictions remained relatively accurate. Field reports documented damaging winds and hail in late June and early July, particularly around Mooresville, Rockport, and Garden City, which caused root lodging and crop damage (Ward, 2023). Because such mechanical damage is not captured by standard soil moisture data, it likely contributed to the yield prediction errors in this region.
Overall, these results indicate that a significant portion of corn yield prediction error is linked to soil moisture prediction error, underscoring the central role of soil moisture in crop growth. By explicitly modeling soil moisture as an intermediate variable, KGML-SM provides a diagnostic layer of interpretability: yield prediction errors can be traced to misrepresentation of soil moisture dynamics. This design not only improves robustness under drought conditions but also offers a transparent explanation for when and why yield predictions fail.
@@
5.5 Strengths, limitations, and future work
Our main contributions are as follows:
  • 1.
    This paper introduced the KGML-SM framework, which integrates process-based and ML models for corn yield prediction while explicitly incorporating the influence of soil moisture.
  • 2.
    A drought-aware loss function was designed to enhance model performance under drought conditions and mitigate overestimation.
  • 3.
While these contributions highlight the novelty and effectiveness of our approach, it is equally important to recognize its limitations and potential directions for future research. First, the study relies on SMAP soil moisture data and APSIM simulations, which are limited in temporal coverage and in the diversity of spatial resolution. Incorporating multi-scale data, such as finer field-level observations and broader regional products, could improve the model’s generalization across heterogeneous environments. Second, the drought-aware loss function is designed to mitigate overestimation under drought conditions, but it does not address other extreme events such as floods or heatwaves. Extending the loss function to include these scenarios would make the framework more comprehensive. Finally, the experiments are limited to the U.S. Corn Belt, and testing the KGML-SM framework in other regions and on different crops will be essential to evaluate its broader applicability. Addressing these limitations will further improve the robustness and generalizability of KGML-SM, ultimately contributing to more reliable and interpretable crop yield prediction under diverse environmental conditions.
@@ 6 Conclusion
In this study, we propose the KGML-SM model, where the W2S encoder is designed to capture the influence of weather on soil moisture, and the attention module is employed to weight different input features for final corn yield prediction. To address the issue that drought conditions often lead to yield overestimation, we introduce a drought-aware loss function to mitigate this problem in drought-affected regions. We construct both an APSIM field-level dataset and a GEE county-level dataset, learning the corn growth process by pretraining KGML-SM on the APSIM field-level dataset and then finetuning it on the GEE county-level dataset. Our analysis covers 12 states in the U.S. Corn Belt to investigate the impact of soil moisture on corn yield prediction. The proposed method consistently outperforms baseline models across multiple test years. Furthermore, we study the spatial and temporal influence of soil moisture through attention visualization, revealing when and where the model places greater focus on soil moisture. Finally, based on the relationship between soil moisture and corn yield prediction, we investigate the causes of prediction inaccuracies and provide explanations. In future work, we aim to apply transfer learning techniques to adapt models trained on well-studied regions with abundant simulated data to regions with limited data availability.